Fermi National Laboratory


Accelerator Update for May 18, 2005 - May 20, 2005

Wednesday May 18
The day shift began with Operations monitoring the Antiprotons Source (Pbar) stack and the Recycler stash, and with MiniBooNE and NuMI taking beam. The Collider Detector at Fermilab (CDF) and D-Zero (D0) experimenters accessed their collision halls.

CDF came out of access at 8:19 AM.

D0 came out of access at 11:19 AM.

Operations put the Tevatron (TeV) into shot setup at 12:11 PM.

At 12:49 PM, the water technicians found a LCW leak in the Meson enclosure MO5 in the MCenter beam line. The techs began their repairs and finished an hour later.

Operations established combination store 4151 at 2:25 PM with an initial luminosity of 117.8E30.

Pbar began stacking at 2:31 PM.

At 3:32 PM, a Booster RF station (BRF16) tripped off and wouldn't reset. Operators bypassed the station, temporarily, and then brought it back up later.

Around 4 PM, Operations established beam to the Meson MCenter and MTest experiments.

Operations put the Recycler into shot setup at 9:02 PM to stash antiprotons.

At 9:46 PM, the TeV suffered a sector E1 quench. The cause was due to an A4 separator spark.

Operations reestablished beam to MCenter at 10:09 PM.

The Recycler stash ended at 10:13 PM.

At 11:20 PM, BRF16 tripped off. It started tripping continually at 11:39 PM and operators bypassed the station at 11:45 PM. Operators contacted an expert.

Thursday May 19
The midnight shift began with Operations monitoring the stash, with Pbar stacking, with operators working on Booster RF problems, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations called in an expert to look at BRF16 at 12:21 AM.

Operations reported at 12:29 AM that TeV experts were tuning in preparation of making orbit changes and TeV studies.

At 1:57 AM, Operation reported that Booster experts temporarily turned off the West Anode Power Supply to begin repairs on BRF16.

Pbar resumed stacking at 2:18 PM.

At 3:24 AM, Operations reported some serious Booster beam problems. Operators investigated and found that BRF15 went down. The Booster expert working on BRF16 thought they might have caused the problem; they will investigate further.

Operations reported a massive lightening storm at 3:58 AM.

TeV studies ended at 4:40 AM.

Operations reported at 5 AM that BRF16 was back online.

Operations put the TeV into shot setup at 5:30 AM.

At 6:25 AM, Operations reported that the Booster notcher tripped off. Operators responded and managed to get it back on by 7 AM.

Operations established store 4154 at 8:43 AM with an initial luminosity of 84.5E30. Pbar resumed stacking and NuMI resumed taking beam.

MCenter resumed taking beam at 8:53 AM.

At 9:18 AM, the TeV suffered a sector A1 quench due to a kicker prefire at A0.

Operations reported at 9:49 AM that they made the CDF and D0 collision halls ready for access. CDF came out of access thirty minutes later.

At 11:28 AM, Operations reported that an Accumulator RF station (ARF1) tripped off. Operators informed experts who investigated the trip. They had it back up by 11:51 AM.

D0 came out of access at 11:52 AM

Operators prepared the TeV for an access at 11:54 AM.

At 1:18 PM, Operations reported that technicians replaced an air solenoid (BV8U) at the TeV sector A4, but that the leak remained. They continued their investigations.

Booster beam returned at 2 PM.

At 2:32 PM, the Tevatron Electron Lens (TEL) tripped off. An operator contacted experts. The experts had in back on line one and a half hours later.

At 2:38 PM, the Booster notcher tripped off. It was reset.

Operations put the TeV through a dry squeeze at 4:18 PM.

Operations established a 4x0 study-store at 8:13 PM.

Operations put the TeV into shot setup at 9:27 PM.

At 10:50 PM, a Booster power supply (MP01) tripped off and wouldn't reset. An operator called in an expert.

At 11:07 PM, BRF16 tripped off. It reset okay.

Friday May 20
The midnight shift began with the TeV in shot setup, with Operations monitoring a small stash, and with MiniBooNE taking beam.

Operations reported at 12:29 AM that MP01 was back online.

At 2:13 AM, Operations reported at sector C4 separator spark.

Operations established store 4157 at 2:24 AM with an initial luminosity of 69.6E30.

Pbar resumed stacking at 2:38 AM.

MCenter resumed taking beam at 2:46 AM.

At 3:33 AM, Operations reported that the TeV suffered a sector E1 quench due to a C4 separator spark.

TeV experts put the separators into conditioning mode at 5:12 AM.

The Plan for Friday and the Weekend
The plan for today is to put the TeV into shot setup as soon as the stack reaches 100mA. The plans for the weekend are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/20/2005   email Fermilab